OpenStudy (anonymous):

rewrite in non fractional form:

OpenStudy (anonymous):

\[\tan ^{2}(x)/(\csc(x)+1)\]

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dfrac{tan ^2x}{cosecx+1}\\ =\dfrac{tan ^2x}{cosecx+1}\times \dfrac{cosecx-1}{cosecx-1}\\ }\end{align}\)

OpenStudy (mathmath333):

rationalize the denominator